4x^2-146x+520=0
(2x- )(2x- )
I can't seem to find the missing #'s
HELP PLEASE!!!

Found 2 solutions by jorel1380, htmentor:
Answer by jorel1380(3719)

4x2-146x+520=0
(2x-65)(2x-8)=0
x=65/2 or 4..

Answer by htmentor(1343)

We can first divide through by 2:
2x^2 - 73x + 260 = 0
The constant term 260 has factors 130 and 2 and 65 and 4.
Recognizing that 73 = 65 + 4*2 and the first factor starts with 2x, we use the last 2 factors:
2x^2 - 73x + 260 = (2x - 65)(x - 4)
Now solve for x:
(2x - 65)(x - 4) = 0
x = 4
2x - 65 = 0 -> x = 65/2 = 32 1/2
Ans: x = 4, 32 1/2
Check using original expression:
4(4)^2 - 146(4) + 520 = 0
4(32.5)^2 - 146(32.5) + 520 = 0